# Hyperparameter Optimization in Machine Learning

### v2

@@ 6.2 Covariance matrix adaptation evolution strategy (CMA-ES)
$\lambda_{k+1}^{j}\sim p_{k}^{j}=\mu_{k}+\sigma_{k}\mathcal{N}(0,\Sigma_{k})\quad\text{for }j\in[N]$
where $N\in\mathbb{N}^{+}$ is the population size, $\mu_{k}$ and $\Sigma_{k}$ are the distribution parameters and $\sigma_{k}>0$ is an overall standard deviation, akin to a step size. Referring to Algorithm 3, this means in CMA-ES $p^{j}_{k}=\mathcal{N}(\mu_{k},\sigma_{k}^{2}\Sigma_{k})$ for all $j\in[N]$.
@@
The covariance matrix is adapted in a similar fashion by reestimating it from the $M$ best points of the previous generation. That is:
$\Sigma_{k}=\sum_{i=1}^{K}w_{i}(\lambda_{k}^{i}-\mu_{k})(\lambda_{k}^{i}-\mu_{k})^{\intercal}$ (6.1)
where the $w_{i}$ are the recombination weights. In practice the update of Eq. (6.1) could be unreliable especially for small populations and hansen2003reducing proposes to maintain a moving average for the covariance matrix. Current implementation of CMA-ES incorporate other strategies to improve robustness and efficiency of the algorithms like adaptation of the step-size $\sigma_{k}$ and generalizations of Eq. (6.1); see [hansen2019pycma].
@@ 6.3 Population-based training (PBT)
Figure 6.1: Evolution of models trained with PBT for Atari game-playing policies. Red dots represent initial points in the parameter and hyperparameter spaces. Paths represents progressions via parameter training. Branching points represent applications of the exploit and explore operations, namely, a copy of the parameters and hyperparamters of better performing individuals onto worse performing one and a subsequent perturbation of their hyperparameter configurations. The performance score is color-coded, with light blue color representing better performing models. Reproduced from [jaderberg2017population].
jaderberg2017population propose a parallel method that maintains a population of model and hyperparameter pairs that are evolved as the models are trained. The procedure, devised in the context of learning deep neural networks in challenging scenarios like image generation or reinforcement learning, consists in interleaving neural network training (by gradient descent) with applications of evolutionary transformations to both the parameter of the models and the hyperparameter of the learning algorithm. PBT is an online hyperparamter optimization algorithm, as hyperparameters are updated alongside model parameters. In other words, in contrast to many techniques we discussed so far, PBT yields an hyperparameter sequence (or schedule) rather than a single hyperparamter configuration. 222 We discuss in more generality hyperparameter schedules and online HPO in Section 8.2.
Suppose one can “decompose” the learning algorithm of interest into a series of stages: $A=A_{T}\circ\dots\circ A_{1}$, where each map $A_{t}$ represents a stage of the learning process that can be warm-started given the outcome of the previous stage. This is particularly straightforward for neural network training: given a space of weights $\mathcal{W}$, each $A_{t}:\mathcal{W}\times\Lambda\to\mathcal{W}$ could be, for instance, an epoch of gradient-based optimization of the network weights (given configuration hyperparameters). Let $N\in\mathbb{N}$ be a population size and $\{(\theta_{0}^{j},\lambda_{0}^{j})\}_{j\in[N]}$ be a set of initial model weights and hyperparameters, where $\theta_{0}^{j}\in\mathcal{W}$ and $\lambda_{0}^{j}\in\Lambda$. These pairs represent the “individuals” in PBT and they can be initialized drawing from suitable distributions.
In PBT, generations correspond to stages of the learning algorithm. At each generation $t\in[T]$ the weights are updated as follows
@@
Then, evolutionary transformations are applied to both parameters and hyperparameters before proceeding to the subsequent learning stage, on the basis of a fitness function $\mathscr{M}$ (e.g. a validation error) evaluated on the current model weights. Suppose we sort individuals in ascending order so that $\mathscr{M}(\tilde{\theta}_{t}^{1})\leq\mathscr{M}(\tilde{\theta}_{t}^{2})\leq\dots\leq\mathscr{M}(\tilde{\theta}_{t}^{N})$. jaderberg2017population propose two type of operations:
  • exploit, a selection strategy that involves copying the parameters and hyperparameters of better performing individuals onto worst performing ones; for instance, one may set $(\theta^{N}_{t},\lambda^{N}_{t})=(\tilde{\theta}^{1}_{t},\lambda^{1}_{t-1})$. jaderberg2017population suggest to replaces the worst-performing 20% with individuals randomly drawn from the best-performing 20%;
  • explore, a mutation transformation that brings over the parameter form the last learning stage, but either resamples the hyperparamter from the starting distribution, or perturbs them, e.g. setting $(\theta_{t}^{j},\lambda_{t}^{j})=(\tilde{\theta}_{t}^{j},\lambda_{t-1}^{j}+\varepsilon)$ for a suitable perturbation $\varepsilon$.
@@
tao2020learning combine PBT with gradient-based optimization of the hyperparameters, which we will discuss in the next chapter.
@@ 6.4 Particle swarm optimization
Particle swarm optimization (PSO) methods draws inspiration from the social behaviour of animals, like ant colonies or bird flocks [kennedy1995particle], as well as from the evolution of particle trajectories in a physical system. The general idea is that each candidate solution evolves according to a simple equation that weights a momentum term, or velocity, and one or more attractor terms to steer the trajectory toward better performing points.
Assume for simplicity that $\Lambda=\mathbb{R}^{m}$, and let $\{(\lambda_{1}^{j},v_{1}^{j})\}_{j=1}^{N}$ be an initial population of hyperparameter and velocity pairs, both drawn randomly from suitable distributions (e.g. uniform distributions). lorenzo2017particle propose an application of PSO to hyperparameter optimization that uses the following equation:
$\lambda_{k+1}^{j}$ $=\lambda_{k}^{j}+v_{k}^{j}$ (6.2)
$v_{k+1}^{j}$ $=\alpha_{1}v_{k}+\alpha_{2}r_{l}\circ(\bar{\lambda}_{k}^{j}-\lambda_{k})+\alpha_{2}r_{g}\circ(\lambda^{*}-\lambda_{k})$ (6.3)
where $\alpha_{1},\alpha_{2},\alpha_{3}\in\mathbb{R}^{+}$ are three coefficients, $r_{l},r_{g}\in\mathbb{R}^{m}$ are two randomly drawn vectors and $\circ$ is the Hadamard product. The configuration $\bar{\lambda}_{k}^{j}$ and $\lambda_{k}^{*}$ are, respectively, the particle’s best and population best points found so far, that is,
@@
where $\bar{\Lambda}^{j}_{k}=\left\{\lambda^{j}_{t}\,:\,t\in[k-1]\right\}$, and
@@
where $\Lambda^{*}_{k}=\left\{\lambda_{t}^{j}\,:\,t\in[k-1],\,j\in[N]\right\}$. The second term in the velocity update of Eq. (6.3) can be interpreted as a local memory attractor that pushes the current configuration toward the individual best configuration found so far, while the third term is a global attractor that pushes toward the global best configuration found so far. The randomness introduced by $r_{l}$ and $r_{g}$ serves to encourage diversity throughout the search. Referring to Algorithm 3, each individual distribution is implicitly modelled by the Equations (6.2)-(6.3).
@@ 6.5 Discussion
Despite their simplicity and their lack of theoretical guarantees, population-based methods enjoy broad applicability and strong empirical performances. In addition, this class of methods is designed to readily scale to parallel and distributed compute systems. Although traditional EA and PSO implementation assume synchronicity, one may easily derive asynchronous versions of many algorithms. For instance in PSO one may replace the global best $\lambda^{*}$ with the best known solution to individual $j$. However, population-based algorithms require a considerable amount of customization due to the presence of several method parameters, task-specific heuristics and design choices, e.g. to define meaningful mutation or perturbation strategies, undermining their accessibility and applicability to novel algorithms and problems. Besides, they typically require a considerable computational power, benefiting usually from large population sizes.
Finally, we note that one may consider successive halving and related methods, which we covered in Chapter 5, as also falling in the category of population-based methods. However, the foundations of those methods are linked to the bandit problem and to hedging strategies [li2017hyperband], rather than taking inspiration from evolutionary principles.
@@
In some cases, the response function, $f(\lambda)$, is differentiable with respect to $\lambda$, enabling the application of optimization algorithms that require the gradients of their objective. Assuming that $f(\lambda)$ be differentiable excludes the direct optimization of discrete hyperparameters such as discrete choices (for example, which nonlinear activation function should be used in a layer of a neural network) or integer-valued hyperparameters (such as the dimension of a layer). There is, however, a distinct advantage of gradient-based methods for HPO, namely the possibility of tuning a (very) high-dimensional hyperparameter vector $\lambda$, which would be rather difficult to do with model-based algorithms. This opportunity allows us to completely rethink when a variable should be considered as a parameter or as a hyperparameter. For example, maclaurin2015gradient suggested to consider the initial weights of a neural network as a large hyperparameter vector, an approach that in their own words effectively “blurs the distinction between learning and meta-learning.” Indeed, gradient-based methods reveal some deep connections between HPO and some forms of meta-learning, which we expand on in Section 8.4). For example, the idea of tuning initial weights from several learning tasks was shortly after successfully exploited in [finn_model-agnostic_2017] in their model-agnostic meta learner.
The crux of gradient-based HPO lies in computing approximate gradients of the response function w.r.t. hyperparameters. In this chapter, we cover three main technical approaches: implicit differentiation (Section 7.2), iterative differentiation (Section 7.3) and fixed-point iterations (Section 7.4). Next, we explore some use cases in Section 7.5. We touch on extensions that allow for the gradient-based optimization of discrete hyperparameters in Section 7.6, and conclude with a discussion centred on computational considerations in Section 7.7.
@@ 7.1 Preliminaries
At a high level, gradient-based HPO methods start from an initial hyperparameter configuration $\lambda_{0}\in\Lambda$ and iteratively and sequentially update such configuration as follows:
@@
where ^$\hat{\nabla}f$ informally denotes some approximation of the gradient of the HPO objective $f$, $\nu>0$ is a step-size and $\mathrm{Proj}_{\Lambda}$ is the projection onto $\Lambda$. Gradient-based optimization of the loss function inside the learning algorithm $A$ is extremely common, especially in deep learning. Hence, to avoid confusion, we will call ^$\hat{\nabla}f$ the hypergradient.
Not every HPO problem is suitable for gradient-based optimization. Recalling the general formulation of the HPO problem of Section 2.2, one underlying assumption shared by all gradient-based HPO methods is that the performance metric $\mathscr{M}$ be differentiable w.r.t. the model output of $A$. In this chapter, to simplify the notation we take the output of $A$ to be a parameter vector in some weight space $\mathcal{W}\subseteq\mathbb{R}^{d}$ and let the performance metric take values in the weight space directly, that is $\mathscr{M}:\mathcal{W}\times\mathscr{D}\ \to\mathbb{R}$. Hence
@@
Hereafter, we will drop training and validation sets.
In order to compute hypergradients, it is also necessary to specify at least some details of the learning algorithm $A(\mathcal{D},\lambda)$ that appears in (2.4) and that we have mostly treated as an opaque box until now. Gradient-based HPO methods are distinguished mainly by the assumptions they make about $A$, which then translate into different computational schemes for ^$\hat{\nabla}f$. Next, we review the main approaches.
@@ 7.2 Hypergradients via implicit differentiation
Many learning algorithms may be conceptually expressed as solutions to optimization problems
@@
where $\mathcal{J}:\mathcal{W}\times\Lambda\to\mathbb{R}$, with $\mathcal{W}\subseteq\mathbb{R}^{d}$ is an objective function (the training error) that depends on the data – we have seen a concrete example in (2.3).  [^1] We leave dependence on data implicit for improving readability. If $\mathcal{J}$ is a smooth function w.r.t. $u$ then, the first-order optimality condition implies that
@@
Notably, neural networks and many related deep learning techniques are formulated in these terms. If $\mathcal{J}$ is twice differentiable w.r.t. $u$ at a point $\bar{\lambda}\in\Lambda$ and $\partial_{w}^{2}\mathcal{J}\in\mathbb{R}^{d\times d}$ is invertible, we can invoke the implicit function theorem [krantz2012implicit] to (7.4) which assures us that (locally) there exist a function $w:U\subset\Lambda\to\mathcal{W}$ such that $\nabla_{u}\mathcal{J}(u,\lambda)_{|u=w(\lambda)}=0$ for all $\lambda\in U$, and $A(\mathcal{D},\lambda)=w(\lambda)$. In such case, we can locally express the HPO objective as the nested function $f(\lambda)=\mathscr{M}(w(\lambda),\mathcal{V})$. Now, if $\mathscr{M}$ is differentiable, we have that
$[\nabla f(\lambda)]^{\intercal}=\partial_{\lambda}f(\lambda)=\partial_{w}\mathscr{M}(w(\lambda),\lambda)\partial_{\lambda}w(\lambda)\qquad\forall\lambda\in U,$ (7.5)
@@
Let us denote by $w_{T}$ an approximate solution of the inner problem, e.g. found with an iterative optimization algorithm. Then, after substituting $w(\lambda)$ with $w_{T}$ and plugging (7.6) into (7.5), one can set $q^{\intercal}=-\partial_{u}\mathscr{M}\left[\partial_{u}^{2}\mathcal{J}\right]^{-1}\in\mathds{R}^{d}$ and approximately solve the linear system
@@
Call $q_{T,K}$ an approximate solution of (7.7). Then an approximate gradient of the outer objective is given by
@@
The two possible sources of approximations stem from ($i$) the solution of the inner problem and ($ii$) of the linear system (7.7). Particular implementation of this computation scheme depend on how (7.7) is solved. One classic procedure is the conjugate gradient (CG) method [hestenes1952methods], which features a linear rate of convergence that depends on the conditioning number of the Hessian of the inner objective.
This computational approach has been studied in [larsen_design_1996, larsen1998adaptive, bengio_gradient-based_2000, chapelle2002choosing, seeger2007cross]. More recently, pedregosa_hyperparameter_2016 proposed an approximation scheme whereby solutions to the inner problems and to the linear equation (7.7) become more precise as the number of HPO iterations $r$ grows, proving convergence for algorithms which feature strongly convex (inner) objectives.

### v3

@@ 6.2 Covariance Matrix Adaptation Evolution Strategy (CMA-ES)
$\lambda_{k+1}^{j}\sim p_{k}^{j}=\mathcal{N}(\mu_{k},\sigma_{k}^{2}\Sigma_{k})\quad\text{for }j\in[N]$
where $N\in\mathds{N}^{+}$ is the population size, $\mu_{k}$ and $\Sigma_{k}$ are the distribution parameters and $\sigma_{k}>0$ is an overall standard deviation, akin to a step size. Referring to Algorithm 3, this means in CMA-ES $p^{j}_{k}=\mathcal{N}(\mu_{k},\sigma_{k}^{2}\Sigma_{k})$ for all $j\in[N]$.
@@
The covariance matrix is adapted similarly by reestimating it from the $M$ best points of the previous generation. That is:
$\Sigma_{k}=\sum_{i=1}^{M}w_{i}(\lambda_{k}^{i}-\mu_{k})(\lambda_{k}^{i}-\mu_{k})^{\intercal}$ (6.1)
where the $w_{i}$ are the recombination weights. In practice the update of Eq. (6.1) could be unreliable especially for small populations and hansen2003reducing proposes to maintain a moving average for the covariance matrix. Current implementation of CMA-ES incorporate other strategies to improve robustness and efficiency of the algorithms like adaptation of the step-size $\sigma_{k}$ and generalizations of Eq. (6.1) [see hansen2019pycma].
@@ 6.3 Population-Based Training (PBT)
Figure 6.1: Evolution of models trained with PBT for Atari game-playing policies. Red dots represent initial points in the parameter and hyperparameter spaces. Paths represents progressions via parameter training. Branching points represent applications of the exploit and explore operations, namely, a copy of the parameters and hyperparameters of better performing individuals onto worse performing one and a subsequent perturbation of their hyperparameter configurations. The performance score is color-coded, with light blue color representing better performing models. Reproduced from [jaderberg2017population].
jaderberg2017population propose a parallel method that maintains a population of model and hyperparameter pairs that are evolved as the models are trained. The procedure, devised in the context of learning deep neural networks in challenging scenarios like image generation or reinforcement learning, consists in interleaving neural network training (by gradient descent) with applications of evolutionary transformations of both the parameters of the models and the hyperparameter of the learning algorithm. PBT is an online hyperparamter optimization algorithm, as hyperparameters are updated alongside model parameters. In other words, in contrast to many techniques we discussed so far, PBT yields a hyperparameter sequence (or schedule) rather than a single hyperparameter configuration.  [^2] We discuss in more generality hyperparameter schedules and online HPO in Section 8.2.
Suppose one can “decompose” the learning algorithm of interest into a series of stages: $A=A_{T}\circ\dots\circ A_{1}$, where each map $A_{t}$ represents a stage of the learning process that can be warm-started given the outcome of the previous stage. This is particularly straightforward for neural network training: given a space of weights $\mathcal{W}$, each $A_{t}\mathrel{\mathop{\ordinarycolon}}\mathcal{W}\times\Lambda\to\mathcal{W}$ could be, for instance, an epoch of gradient-based optimization of the network weights. Let $N\in\mathds{N}$ be a population size and $\{(\theta_{0}^{j},\lambda_{0}^{j})\}_{j\in[N]}$ be a set of initial model weights and hyperparameters, where $\theta_{0}^{j}\in\mathcal{W}$ and $\lambda_{0}^{j}\in\Lambda$. These pairs represent the “individuals” in PBT and they can be initialized drawing from suitable distributions.
Generations in PBT correspond to stages of the learning algorithm. At each generation $t\in[T]$ the weights are updated as follows:
@@
Then, evolutionary transformations are applied to both parameters and hyperparameters before proceeding to the subsequent learning stage, on the basis of a fitness function $\mathscr{M}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to\mathds{R}$ (e.g. a validation error, cf. Section 2.2) evaluated on the current model weights.  [^3] Note that here we cannot explicitly refer to the response function $f$, as hyperparameters may change throughout iterations; see also the discussion on hyperparamter schedules in Section 8.2. Suppose we sort individuals in ascending order so that $\mathscr{M}(\tilde{\theta}_{t}^{1})\leq\mathscr{M}(\tilde{\theta}_{t}^{2})\leq\dots\leq\mathscr{M}(\tilde{\theta}_{t}^{N})$. jaderberg2017population propose two types of operations:
  • Exploit: a selection strategy that involves copying the parameters and hyperparameters of better performing individuals onto worst performing ones; for instance, one may set $(\theta^{N}_{t},\lambda^{N}_{t})=(\tilde{\theta}^{1}_{t},\lambda^{1}_{t-1})$. jaderberg2017population suggest to replace the worst-performing 20% with individuals randomly drawn from the best-performing 20%;
  • Explore: a mutation transformation that brings over the parameter form the last learning stage, but either resamples the hyperparamters from the starting distribution, or perturbs them, e.g. by setting $(\theta_{t}^{j},\lambda_{t}^{j})=(\tilde{\theta}_{t}^{j},\lambda_{t-1}^{j}+\varepsilon)$ for a perturbation $\varepsilon$ such that $\lambda_{t-1}^{j}+\varepsilon\in\Lambda$.
@@
tao2020learning combine PBT with gradient-based optimization of the hyperparameters. The authors propose an “hyperparameter mutation” algorithm that learns a dynamic trade-off between local search via gradient-based hyperparameter optimization and global exploration via PBT. We discussed gradient-based HPO methods in the next chapter.
@@ 6.4 Particle Swarm Optimization
Particle swarm optimization (PSO) methods draw inspiration from the social behavior of animals, like ant colonies or bird flocks [kennedy1995particle], as well as from the evolution of particle trajectories in a physical system. The general idea is that each candidate solution evolves according to a simple equation that weights a momentum term, or velocity, and one or more attractor terms to steer the trajectory toward better performing points.
Assume for simplicity that $\Lambda=\mathds{R}^{m}$, and let $\{(\lambda_{1}^{j},v_{1}^{j})\}_{j=1}^{N}$ be an initial population of hyperparameter and velocity pairs, both drawn randomly from suitable distributions (e.g., uniform distributions). lorenzo2017particle propose an application of PSO to hyperparameter optimization that uses the following equation:
$\lambda_{k+1}^{j}$ $=\lambda_{k}^{j}+v_{k}^{j},$ (6.2)
$v_{k+1}^{j}$ $=\alpha_{1}v_{k}+\alpha_{2}r_{l}\circ(\bar{\lambda}_{k}^{j}-\lambda_{k})+\alpha_{2}r_{g}\circ(\lambda^{*}-\lambda_{k}),$ (6.3)
where $\alpha_{1},\alpha_{2},\alpha_{3}\in\mathds{R}^{+}$ are three coefficients, $r_{l},r_{g}\in\mathds{R}^{m}$ are two randomly drawn vectors and $\circ$ is the Hadamard product. The configuration $\bar{\lambda}_{k}^{j}$ and $\lambda_{k}^{*}$ are, respectively, the best particle and population points found so far, that is,
@@
where $\bar{\Lambda}^{j}_{k}=\left\{\lambda^{j}_{t}\in\Lambda\mid t\in[k-1]\right\}$, and
@@
where $\Lambda^{*}_{k}=\left\{\lambda_{t}^{j}\in\Lambda\mid t\in[k-1],\,j\in[N]\right\}$. The second term in the velocity update of Eq. (6.3) can be interpreted as a local memory attractor that pushes the current configuration toward the individual best configuration found so far, while the third term is a global attractor that pushes toward the global best configuration found so far. The randomness introduced by $r_{l}$ and $r_{g}$ serves to encourage diversity throughout the search. Referring to Algorithm 3, each individual distribution is implicitly modeled by the Eqs. (6.2)-(6.3).
@@ 6.5 Discussion
Despite the lack of theoretical guarantees, population-based methods enjoy broad applicability and strong empirical performances. In addition, this class of methods is designed to readily scale to parallel and distributed compute systems. Although traditional EA and PSO implementation assume synchronicity, one may easily derive asynchronous versions of many algorithms. For instance in PSO, one may replace the global best configuration $\lambda^{*}$ with the best known solution to each individual $j$. However, population-based algorithms require a considerable amount of customization due to the presence of several method parameters, task-specific heuristics and design choices. For instance, it might be challenging to define meaningful mutation or perturbation strategies, or implement effective niching. These considerations may undermine the applicability of these methods to novel learning algorithms and HPO problems, and may preclude accessibility to non-expert users. Besides, they typically require a considerable computational power, benefiting usually from large population sizes.
@@
In some cases, the response function, $f(\lambda)$, is differentiable with respect to $\lambda$, enabling the application of optimization algorithms that require the gradients of their objective. Assuming $f(\lambda)$ to be differentiable excludes the direct optimization of discrete hyperparameters such as discrete choices, for example, which nonlinear activation function should be used in a layer of a neural network, or integer-valued hyperparameters, such as the dimension of a layer. There is, however, a distinct advantage of gradient-based methods for HPO, namely the possibility of tuning a (very) high-dimensional hyperparameter vector $\lambda$, which would be rather difficult to do with model-based algorithms. This opportunity allows us to completely rethink when a variable should be considered as a parameter or as a hyperparameter. maclaurin2015gradient suggested considering the initial weights of a neural network as a large hyperparameter vector, an approach that in their own words effectively “blurs the distinction between learning and meta-learning.” Indeed, gradient-based methods reveal some deep connections between HPO and some forms of meta-learning, which we expand on in Section 8.4. For example, the idea of tuning initial weights from several learning tasks was successfully exploited by finn_model-agnostic_2017 in their model-agnostic meta learner.
The crux of gradient-based HPO lies in computing approximate gradients of the response function w.r.t. hyperparameters. In this chapter, we cover three main technical approaches: implicit differentiation (Section 7.2), iterative differentiation (Section 7.3) and fixed-point iterations (Section 7.4). Next, we explore some use cases in Section 7.5. We touch on extensions that allow for the gradient-based optimization of discrete hyperparameters in Section 7.5.1, and conclude with a discussion centered on computational considerations in Section 7.6.
@@ 7.1 Preliminaries
At a high level, gradient-based HPO methods start from an initial hyperparameter configuration $\lambda_{0}\in\Lambda$ and iteratively update such configuration as follows:
@@
where $\hat{\nabla}f$ informally denotes some approximation of the gradient of the HPO objective $f$, $\nu>0$ is a step-size and $\mathrm{Proj}_{\Lambda}$ is the projection onto $\Lambda$. Gradient-based optimization of the loss function inside the learning algorithm $A$ is extremely common, especially in deep learning. Hence, to avoid confusion, we will call $\hat{\nabla}f$ the hypergradient.
Not every HPO problem is suitable for gradient-based optimization. Recalling the general formulation of the HPO problem in Section 2.2, one underlying assumption shared by all gradient-based HPO methods is that the performance metric $\mathscr{M}$ be differentiable w.r.t. the model output of $A$. In this chapter, to simplify the notation we take the output of $A$ to be a parameter vector $w\in\mathcal{W}\subseteq\mathbb{R}^{d}$ and let the performance metric take values in the weight space directly, that is $\mathscr{M}\mathrel{\mathop{\ordinarycolon}}\mathcal{W}\times\mathscr{D}\ \to\mathbb{R}$. Hence,
@@
In the following, it will be useful to explicitly write $w$ as a function of $\lambda$, but we will drop $\mathcal{D}$ and $\mathcal{V}$ to simplify the notation.
In order to compute hypergradients, it is also necessary to specify at least some details of the learning algorithm $A(\mathcal{D},\lambda)$ that appears in (2.4) and that we have mostly treated as an opaque box until now. Gradient-based HPO methods are distinguished mainly by the assumptions they make about $A$, which then translate into different computational schemes for $\hat{\nabla}f$. Next, we review the main approaches.
@@ 7.2 Hypergradients via Implicit Differentiation
Many learning algorithms may be conceptually expressed as solutions to optimization problems:
@@
where $\mathcal{J}\mathrel{\mathop{\ordinarycolon}}\mathcal{W}\times\Lambda\to\mathbb{R}$ is an objective function (e.g., the training loss) that depends on the data, and $\mathcal{W}\subseteq\mathbb{R}^{d}$ is a parameter vector. We have seen a concrete example in Eq. (2.3). If $\mathcal{J}$ is a smooth function w.r.t. $u$ then, the first-order optimality condition implies that
@@
Notably, neural networks and many related deep learning techniques are formulated in these terms. If $\mathcal{J}$ is twice differentiable w.r.t. $u$ at a point $\bar{\lambda}\in\Lambda$ and $\partial_{w}^{2}\mathcal{J}(u,\lambda)\in\mathbb{R}^{d\times d}$ is invertible, we can invoke the implicit function theorem [see, e.g., krantz2012implicit] to Eq. (7.4) which assures us that (locally) there exist a function $w\mathrel{\mathop{\ordinarycolon}}U\subset\Lambda\to\mathcal{W}$ such that $\nabla_{u}\mathcal{J}(u,\lambda)_{|u=w(\lambda)}=0$ for all $\lambda\in U$, and $A(\mathcal{D},\lambda)=w(\lambda)$. We can then locally express the HPO objective as the nested function $f(\lambda)=\mathscr{M}(w(\lambda),\mathcal{V})$. Now, if $\mathscr{M}$ is differentiable, we have that
$\nabla f(\lambda)]^{\intercal}=\partial_{\lambda}f(\lambda)=\partial_{w}\mathscr{M}\left(w(\lambda),\mathcal{V}\right)\partial_{\lambda}w(\lambda)\qquad\forall\lambda\in U,$ (7.5)
@@
Let us denote by $w_{T}$ an approximate solution of the inner problem, e.g. found with an iterative optimization algorithm. Then, after substituting $w(\lambda)$ with $w_{T}$ and plugging Eq. (7.6) into Eq. (7.5), one can set $q^{\intercal}=-\partial_{u}\mathscr{M}\left[\partial_{u}^{2}\mathcal{J}\right]^{-1}\in\mathbb{R}^{d}$ and approximately solve the linear system
@@
Call $q_{T,K}$ an approximate solution of Eq. (7.7). Then an approximate gradient of the outer objective is given by
@@
The two possible sources of approximations stem from ($i$) the solution of the inner problem and ($ii$) of the linear system in Eq. (7.7). A particular implementation of this computation scheme depends on how Eq. (7.7) is solved. One classic procedure is the conjugate gradient (CG) method [hestenes1952methods], which features a linear rate of convergence that depends on the conditioning number of the Hessian of the inner objective.
This computational approach has been studied by larsen_design_1996, larsen1998adaptive, bengio_gradient-based_2000, chapelle2002choosing, seeger2007cross. More recently, pedregosa_hyperparameter_2016 proposed an approximation scheme whereby solutions to the inner problems and to the linear Eq. (7.7) become more precise as the number of HPO iterations $r$ grows, proving convergence for algorithms which feature strongly convex (inner) objectives.